A company is going to install a new automated plastic molding press that will increase

The expected annual income.

Two different presses are available.

The initial cost, annual expense, and the expected increase in annual income for each

press is shown for a five year period. There is no expected salvage value for either press.

Based on an economic study using an interest rate of 10% which press should be chosen?

                                                                Machine A                                           Machine B     

Initial cost                                           $128,000                                              $142,000

Annual expenses                             $54,500                                                 $48,600

Increased annual income              $115,700                                              $114,500
Answer #1

NPW of A = -128000 + (115700 - 54500)*(P/A,10%,5)

= -128000 + (115700 - 54500)*3.790787

= 103996.16

NPW of B = -142000 + (114500 - 48600)*(P/A,10%,5)

= -142000 + (114500 - 48600)*3.790787

= 107812.86

as NPW of B is more, it should be selected